OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 3?

OpenStudy (welshfella):

common difference = second term - first term

OpenStudy (welshfella):

explicit equation is an = a + (n - 1)d where common difference = d and first term = a

OpenStudy (anonymous):

I think it starts like a_n=5-2(n-1) right?

OpenStudy (welshfella):

yes good job

OpenStudy (welshfella):

now the domain is the possible value of the terms

OpenStudy (welshfella):

listing the terms as the value odf d = -2 they are 5, 3, 1, -1, - 3......

OpenStudy (anonymous):

the last part I need is whether or not it is all integers where n>0 or n>1

OpenStudy (welshfella):

n starts from 1 so n > 0

OpenStudy (anonymous):

alright thank you again!

OpenStudy (welshfella):

yw